# ReCLIP++: Learn to Rectify the Bias of CLIP for Unsupervised Semantic Segmentation

### v1

@@ Abstract
Recent works utilize CLIP to perform the challenging unsupervised semantic segmentation task where only images without annotations are available. However, we observe that when adopting CLIP to such a pixel-level understanding task, unexpected bias (including class-preference bias and space-preference bias) occurs. Previous works don’t explicitly model the bias, which largely constrains the segmentation performance. In this paper, we propose to explicitly model and rectify the bias existing in CLIP to facilitate the unsupervised semantic segmentation task. Specifically, we design a learnable “Reference” prompt to encode class-preference bias and a projection of the positional embedding in vision transformer to encode space-preference bias respectively. To avoid interference, two kinds of biases are firstly independently encoded into different features, i.e., the Reference feature and the positional feature. Via a matrix multiplication between Reference feature and positional feature, a bias logit map is generated to explicitly represent two kinds of biases. Then we rectify the logits of CLIP via a simple element-wise subtraction. To make the rectified results smoother and more contextual, we design a mask decoder which takes the feature of CLIP and rectified logits as input and outputs a rectified segmentation mask with the help of Gumbel-Softmax operation. To make the bias modeling and rectification process meaningful and effective, a contrastive loss based on masked visual features and the text features of different classes is imposed. To further improve the segmentation, we distill the knowledge from the rectified CLIP to the advanced segmentation architecture via minimizing our designed mask-guided, feature-guided and text-guided loss terms. Extensive experiments on various benchmarks including PASCAL VOC, PASCAL Context, ADE20K, Cityscapes and COCO Stuff demonstrate that our method performs favorably against previous state-of-the-arts. The implementation is available at: https://github.com/dogehhh/ReCLIP.
Index Terms:
@@
Unsupervised semantic segmentation, contrastive loss, CLIP, bias rectification.
I Introduction
Figure 1: (a) Space-preference bias. (Left): The relationship between distance ($x$-axis) and mIoU ($y$-axis) is drawn on PASCAL VOC [1]. The distance means the spatial distance between the object centriod and the image centriod and mIoU is computed based on predictions and ground truth. The curve shows that CLIP [2] (green) is apparently better at segmentation for central objects than boundary ones, but our method (blue) effectively mitigates this bias. More details about how we draw this figure has been shown in our appendix. (Right): Visualizations also show our improvement on space-preference bias qualitatively. (b) Class-preference bias. (Left): We randomly select 6 classes from PASCAL Context [3] and draw the confusion matrix of CLIP and our model. It shows that besides the ground truth, CLIP also prefers to assign an incorrect but relevant label to a pixel in quite a few cases, while our results show apparent improvement. (Right): The visualizations are consistent with what we observed in confusion matrix. For example, for a “cow”, CLIP tends to classify it as “dog” incorrectly.
@@ I Introduction
Semantic segmentation aims to attach a semantic label to each pixel of an image. Since the rising of deep learning [4, 5, 2, 6, 7], semantic segmentation has been widely adopted in real-world applications, e.g., autonomous driving, medical image segmentation, etc. Conventional approaches [8, 9, 10, 11, 12, 13, 14] for semantic segmentation have achieved remarkable performance. However, the superior performance of those methods relies heavily on large amounts of fully annotated masks. Collecting such high-quality pixel-level annotations can be both time-consuming and expensive, e.g., some annotations for specialized tasks require massive expert knowledge, some are even inaccessible due to privacy reasons, etc. Therefore, it is necessary to explore unsupervised semantic segmentation where only images without annotations are available.
@@
Unsupervised semantic segmentation (USS) has been studied for years. Many non-language-guided USS methods have been proposed, e.g., clustering-based methods [15, 16, 17, 18, 19], contrastive-learning-based methods [20, 21] and boundary-based methods [22], etc. Despite promising progress achieved, there still exhibits a large performance gap between USS and the supervised segmentation methods. Besides, these methods typically obtain class-agnostic masks and have to depend on additional processing (e.g., Hungarian matching) to assign semantic labels to the masks, rendering them less practical in real scenarios.
@@
Recently, large-scale visual-language pre-trained models, e.g., CLIP [2], demonstrate superior zero-shot classification performance by comparing the degree of alignment between the image feature and text features of different categories. A few CLIP-based USS approaches [23, 24, 25, 26, 27] emerge and show remarkable performance improvement compared with the non-language-guided USS methods. These models require no access to any types of annotations, and directly assign a label to each pixel, benefiting from the aligned vision and text embedding space of CLIP. However, good alignment between image-level visual feature and textual feature doesn’t necessarily mean good alignment between pixel-level visual feature and textual feature. Thus, for CLIP, unexpected bias may inevitably appear. Previous works don’t explicitly model the bias, which may largely constrain their segmentation performance.
@@
We observe two kinds of biases existing in CLIP. From one aspect, as shown in Fig. 1 (a), CLIP exhibits space-preference bias. CLIP performs apparently better for segmenting central objects than the ones distributed near the image boundary. It can be reflected by the fact that mIoU decreases as the distance between the object centroid and the image centroid increases. From the other aspect, as shown in Fig. 1 (b), there exists class-preference bias between relevant categories in CLIP. For example, according to visualizations (right), when ground truth is “cow”, CLIP tends to incorrectly classify it as “dog”. We also show such trend between randomly selected classes by confusion matrix (left). Elements on diagonal line represent right classification, while others are false. We observe that besides ground truth, CLIP usually prefers to assign an incorrect but relevant label to a pixel in quite a few cases, exhibiting a wide range of class-preference bias of CLIP.
@@
In this paper, we propose to explicitly model and rectify the bias of CLIP to facilitate the USS. Specifically, we design two kinds of text inputs for each class, which are named as “Query” and “Reference” respectively. The Query is manually designed and fixed while the Reference contains learnable prompts. We pass the Query and Reference through the text encoder of CLIP to obtain Query feature and Reference feature. Then we adopt Query feature as the segmentation head to generate the Query logits for each pixel of the image, which represents the segmentation ability of original CLIP. In contrast, we utilize Reference feature to encode the class-preference bias. In addition, we project the positional embedding of CLIP’s vision transformer to generate positional feature, which encodes the space-preference bias. The encoding processes of two biases are independent to avoid interference. Via a matrix multiplication between Reference feature and positional feature, a bias logit map is generated to explicitly represent two kinds of biases. Then, we remove the bias from original CLIP via a logit-subtraction mechanism, i.e., subtracting the bias logits from the Query logits. To make the rectified results smoother and more contextual, we concatenate the rectified logit map to the feature extracted by CLIP’s visual encoder and pass them through a designed decoder. Then we apply the Gumbel-Softmax operation to the output of decoder to generate a rectified mask. To make the bias modeling and rectification process meaningful and effective, the contrastive loss based on masked visual features (i.e., applying rectified masks to the visual feature of CLIP and pooling) and text features of different categories is imposed. Finally, we distill the knowledge from the rectified CLIP to the advanced segmentation architecture to further improve the segmentation performance, via our designed mask-guided, feature-guided and text-guided distillation loss terms.
@@
We conduct extensive experiments on five standard semantic segmentation benchmarks, including PASCAL VOC [1], PASCAL Context [3], ADE20K [28], Cityscapes [29] and COCO Stuff [30]. Experiment results demonstrate that ReCLIP++ performs favourably against previous state-of-the-arts. Notably, on PASCAL VOC, our method outperforms MaskCLIP+ [23] by 9.2% and CLIP-S4 [26] by 7.2% mIoU. Extensive ablation studies verify the effectiveness of each design in our framework.
@@
This paper extends our conference version [27]. To differentiate, we denote the method proposed in our conference version as ReCLIP and that in this paper as ReCLIP++. The key technical improvements of ReCLIP++ beyond ReCLIP are summarized as follows: 1) We optimize the design of Bias Extraction Module. Specifically, we independently encode class-preference bias and space-preference bias into class-level Reference feature and patch-level positional feature respectively. Then the bias logit map, which explicitly represents both biases, is obtained by combining two features with matrix multiplication. 2) We additionally introduce a mask decoder, which takes the rectified logit map and the visual feature of CLIP as input and outputs smoother and more contextual rectified predictions. 3) We design a new strategy to generate more accurate multi-label hypothesis for each image, which provides a better supervision for the bias rectification process. 4) We evaluate our methods on two more datasets including Cityscapes [29] and COCO Stuff [30]. 5) The segmentation performance of ReCLIP++ outperforms ReCLIP remarkably on all the datasets.
@@
  • We observe that when applying CLIP to pixel-level understanding tasks, unexpected bias including space-preference bias and class-preference bias, occurs. The bias may largely constrain the segmentation performance of CLIP-based segmentation models.
@@
  • We propose to explicitly encode the class-preference and space-preference bias of CLIP via learnable Reference text inputs and projection of positional embedding respectively, and model both biases into one bias logit map via a matrix multiplication. Through a simple logit-subtraction mechanism and the contrastive loss built on masked features of different classes, we effectively rectify the bias of CLIP.
@@
  • We conduct extensive experiments on segmentation benchmarks under the unsupervised setting. Experiment results show superior performance of our method to previous state-of-the-arts.
II Related Work
@@
Pre-trained vision-language models. Pre-trained vision-language models (VLMs) [31, 32, 33, 34, 35] have developed rapidly with the help of large-scale image-text pairs available on the Internet. Recently, CLIP [2], ALIGN [36] and Slip [37] have made great progress on learning visual and textual representations jointly by using contrastive learning. With the image-level alignment with text, pre-trained VLMs have strong ability for zero-shot classification task and can be transferred to various downstream tasks, such as object detection [38, 39] and semantic segmentation [23].
@@
Unsupervised semantic segmentation. While conventional approaches of semantic segmentation [8, 9, 11, 10, 14] rely on pixel-level annotations and weakly-supervised methods [40, 41, 42, 43] still ask for image-level labels, unsupervised semantic segmentation (USS) methods [16, 44, 18, 20, 45] explore to train a segmentation model without any annotations. Models like [46, 47] adopt generative model [48] to separate foreground with background or generate corresponding masks. SegSort [22], HSG[17] and ACSeg [19] use clustering strategy, while IIC [15] uses mutual information maximization to perform unsupervised learning. MaskContrast [20] introduces contrastive learning into USS. Others like DSM [49] and LNE [50] exploit features extracted from self-supervised model and spectral graph theory to facilitate segmentation. However, the methods mentioned above either fail to segment images with multi-category objects or show a large performance gap with the supervised methods. Besides, they can only obtain class-agnostic masks and have to depend on additional strategy, such as Hungarian-matching algorithm [51], to match corresponding category with the segment. Recently, pre-trained vision-language models are adopted in USS. MaskCLIP [23] modifies the image encoder of CLIP to generate patch-level features and directly performs segmentation with text features as classifiers. CLIP-py [25] performs contrastive learning between visual features from self-supervised ViT [52] and text features from CLIP. ReCo [24] performs image retrieval with CLIP and extracts class-wise embedding as classifier with co-segmentation. CLIP-S4 [26] learns pixel embeddings with pixel-segment contrastive learning and aligns such embeddings with CLIP in terms of embedding and semantic consistency. These methods can directly assign a label to each pixel, which falls into the category of language-guided unsupervised semantic segmentation. However, directly applying CLIP may result in unexpected bias. No previous method considers such bias. In this paper, we propose to explicitly model and rectify the bias of CLIP for unsupervised semantic segmentation.
@@
Language-guided semantic segmentation. Recently, many works explore semantic segmentation guided by language under different settings. Zero-shot works [53, 54, 55, 56] split classes into seen and unseen set. During the training period, only masks of seen classes are provided. For inference, models are tested on both seen and unseen classes, but the test data is still in the same distribution with the training data. Open-vocabulary works [57, 58, 59, 60, 61, 62, 63] are trained in one scenario with extra annotations including class-agnostic masks or image captions, but are used for predicating segmentation masks of novel classes in other scenarios. From the technical view, our method also falls into the category of language-guided semantic segmentation. However, we consider the unsupervised setting. In this setting, we have access to images without any annotations during training. The training and inference images are sampled from the same distributions and the same set of categories. Such a setting is different to the typical zero-shot or open-vocabulary setting.
III Method
@@
Overview We aim to rectify the bias of CLIP including the class-preference bias and the space-preference bias, to facilitate unsupervised semantic segmentation. From a high level, class-preference bias reflects the shift of CLIP predictions towards specific classes, while space-preference bias reflects the shift of CLIP predictions towards specific spatial locations. Both biases will be finally reflected in the bias logit map. A reasonable way to rectify the bias is to subtract the bias logit map from the normal logit map predicted by original CLIP.
@@
The general framework of our method is illustrated in Fig. 2, consisting of a rectification stage and a distillation stage. At the rectification stage (see Fig. 2 (a)), we first forward the image $I\in\mathbb{R}^{3\times H\times W}$ through the image encoder of CLIP to obtain the patch-level image feature $Z$. For each class, we manually design the text input which is named Query $Q$ and fix $Q$ throughout the training. We pass $Q$ through the text encoder of CLIP to obtain query text feature $W_{q}$ for each class, which serves as the weight of query segmentation head. With $W_{q}$, we can obtain a query logit map $M_{q}$, which represents the segmentation ability of original CLIP (Sec. III-A).
@@
We extract the bias existing in CLIP in a learnable way. Specifically, we design a kind of learnable text input named Reference $R$ for each class. Passing $R$ through the text encoder of CLIP, we obtain Reference text feature $W_{r}$, which is expected to encode the class-preference bias. Meanwhile, the positional embedding is projected into the positional feature $W_{p}$ to encode the space-preference bias. Then the bias logit map $M_{b}$ is extracted via a matrix multiplication between $W_{r}$ and $W_{p}$ (Sec. III-B).
@@
The Query logit map and the bias logti map are then passed through a Rectified Mask Generation module to generate rectified semantic masks. Firstly, a rectified logit map $M$ is generated via a simple subtraction operation between $M_{q}$ and $M_{b}$. Then, a mask decoder takes $M$ and visual feature $Z$ as input and outputs smoother and more contextual semantic masks with the help of Gumbel-Softmax operation (Sec. III-C).
@@
To enable a more meaningful and effective bias rectification, we generate multi-label hypothesis for each image and impose contrastive loss based on masked visual features and Query text features of different classes (Sec. III-D).
At the distillation stage (see Fig. 2 (b)), we distill the knowledge of the rectified CLIP to the advanced segmentation architecture with specifically designed mask-guided, feature-guided and text-guided loss terms (Sec. III-E). In both the rectification and distillation stages, CLIP is kept frozen.
@@
Figure 2: Method overview of ReCLIP++. We propose a new framework for language-guided unsupervised semantic segmentation. (a) Rectification Stage: At this stage, we aim to rectify the class-preference and space-preference bias of CLIP via specifically designed Bias Extraction module, Rectified Mask Generation module and Contrastive Loss module. (b) Distillation Stage: We distill knowledge from the rectified CLIP to the advanced segmentation architecture with the mask-guided, feature-guided and text-guided loss terms. The “Rectified Mask Generation” module and the “Contrastive Loss” module in the figure is same to those in the Rectification Stage respectively.
III-A Baseline: Directly Segment with CLIP

### v2

@@ Abstract
Recent works utilize CLIP to perform the challenging unsupervised semantic segmentation task where only images without annotations are available. However, we observe that when adopting CLIP to such a pixel-level understanding task, unexpected bias (including class-preference bias and space-preference bias) occurs. Previous works don’t explicitly model the bias, which largely constrains the segmentation performance. In this paper, we propose to explicitly model and rectify the bias existing in CLIP to facilitate the unsupervised semantic segmentation task. Specifically, we design a learnable “Reference” prompt to encode class-preference bias and a projection of the positional embedding in the vision transformer to encode space-preference bias respectively. To avoid interference, two kinds of biases are firstly independently encoded into different features, i.e., the Reference feature and the positional feature. Via a matrix multiplication between the Reference feature and the positional feature, a bias logit map is generated to explicitly represent two kinds of biases. Then we rectify the logits of CLIP via a simple element-wise subtraction. To make the rectified results smoother and more contextual, we design a mask decoder which takes the feature of CLIP and the rectified logits as input and outputs a rectified segmentation mask with the help of Gumbel-Softmax operation. A contrastive loss based on the masked visual features and the text features of different classes is imposed, which makes the bias modeling and rectification process meaningful and effective. Extensive experiments on various benchmarks including PASCAL VOC, PASCAL Context, ADE20K, Cityscapes, and COCO Stuff demonstrate that our method performs favorably against previous state-of-the-arts. The implementation is available at: https://github.com/dogehhh/ReCLIP.
keywords:
@@
Unsupervised semantic segmentation, CLIP, Bias rectification, Contrastive learning
1 Introduction
@@
Semantic segmentation aims to attach a semantic label to each pixel of an image. Since the rise of deep learning [1, 2, 3, 4, 5], semantic segmentation has been widely adopted in real-world applications, e.g., autonomous driving, medical image segmentation, etc. Conventional approaches [6, 7, 8, 9] for semantic segmentation have achieved remarkable performance. However, the superior performance of those methods relies heavily on large amounts of fully annotated masks. Collecting such high-quality pixel-level annotations can be both time-consuming and expensive, e.g., some annotations for specialized tasks require massive expert knowledge, some are even inaccessible due to privacy reasons, etc. Therefore, it is necessary to explore unsupervised semantic segmentation where only images without annotations are available.
@@
Unsupervised semantic segmentation (USS) has been studied for years. Many non-language-guided USS methods have been proposed, e.g., clustering-based methods [10, 11, 12, 13, 14, 15], contrastive-learning-based methods [16, 17] and boundary-based methods [18], etc. Despite promising progress achieved, there still exhibits a large performance gap between USS and the supervised segmentation methods. Besides, these methods typically obtain class-agnostic masks and have to depend on additional processing (e.g., Hungarian matching) to assign semantic labels to the masks, rendering them less practical in real scenarios.
@@
Recently, large-scale visual-language pre-trained models, e.g., CLIP [19], demonstrate superior zero-shot classification performance by comparing the degree of alignment between the image feature and text features of different categories. A few CLIP-based USS approaches [20, 21, 22, 23, 24] emerge and show remarkable performance improvement compared with the non-language-guided USS methods. These models require no access to any types of annotations, and can directly assign a label to each pixel, benefiting from the aligned vision and text embedding space of CLIP. However, good alignment between image-level visual feature and textual feature doesn’t necessarily mean good alignment between pixel-level visual feature and textual feature. Thus, for CLIP, unexpected bias may inevitably appear. Previous works don’t explicitly model such bias, which may largely constrain their segmentation performance.
Figure 1: (a) Space-preference bias. (Left): The relationship between distance ($x$-axis) and mIoU ($y$-axis) is drawn on PASCAL VOC [25]. The distance means the spatial distance between the object centroid and the image centroid and mIoU is computed based on predictions and ground truth. The curve shows that MaskCLIP [20] (pink) is apparently better at segmentation for central objects than boundary ones, but our method ReCLIP++ (blue) effectively mitigates this bias. More details about how we draw this figure have been shown in our appendix. (Right): Visualizations also show our improvement in space-preference bias qualitatively. (b) Class-preference bias. (Left): We randomly select 6 classes from PASCAL Context [26] and draw the confusion matrix of MaskCLIP and our model. It shows that besides the ground truth, MaskCLIP also prefers to assign an incorrect but semantically relevant label to a pixel in quite a few cases, while our results show an apparent improvement. (Right): The visualizations are consistent with what we observed in the confusion matrix. For example, for a “sheep” (dark yellow), MaskCLIP tends to classify it as a “cow” (green) incorrectly.
@@
When directly adopting CLIP in USS like MaskCLIP [20], we observe two kinds of biases existing in CLIP. From one aspect, as shown in Fig. 1 (a), CLIP exhibits space-preference bias. CLIP performs apparently better for segmenting central objects than the ones distributed near the image boundary. It can be reflected by the fact that mIoU decreases as the distance between the object centroid and the image centroid increases. From the other aspect, as shown in Fig. 1 (b), there exists class-preference bias between semantically relevant categories in CLIP. For example, according to visualizations (right), when the ground truth is a “sheep” (dark yellow), CLIP tends to incorrectly classify it as a “cow” (green). We also show such a trend between randomly selected classes by confusion matrix (left). Elements on the diagonal line represent the right classification, while others are false. We observe that besides ground truth, CLIP usually prefers to assign an incorrect but semantically relevant label to a pixel in quite a few cases, exhibiting a wide range of class-preference bias of CLIP.
@@
In this paper, we propose to explicitly model and rectify the bias of CLIP to facilitate the USS. Specifically, we design two kinds of text inputs for each class, which are named “Query” and “Reference” respectively. The Query is manually designed and fixed while the Reference contains learnable prompts. We pass the Query and Reference through the text encoder of CLIP to obtain the Query feature and Reference feature. We adopt the Query feature as the segmentation head to generate the Query logits for each pixel of the image, which represents the original segmentation ability of vanilla CLIP. In contrast, we utilize the Reference feature to encode the class-preference bias. In addition, we project the positional embedding of CLIP’s vision transformer to generate positional feature, which encodes the space-preference bias. The encoding processes of two biases are independent to avoid interference. Via a matrix multiplication between the Reference feature and the positional feature, a bias logit map is generated to explicitly represent two kinds of biases. Then, we remove the bias from the original CLIP via a logit-subtraction mechanism, i.e., subtracting the bias logits from the Query logits. To make the rectified results smoother and more contextual, we concatenate the rectified logit map to the feature extracted by CLIP’s visual encoder and pass them through a designed decoder. Then we apply the Gumbel-Softmax operation to the output of the decoder to generate a rectified mask. To make the bias modeling and rectification process meaningful and effective, the contrastive loss based on masked visual features (i.e., applying rectified masks to the visual feature of CLIP and pooling) and text features of different categories is imposed.
@@
We conduct extensive experiments on five standard semantic segmentation benchmarks, including PASCAL VOC [25], PASCAL Context [26], ADE20K [27], Cityscapes [28] and COCO Stuff [29]. Experiment results demonstrate that ReCLIP++ performs favorably against previous state-of-the-arts. Notably, on PASCAL VOC, our method outperforms MaskCLIP+ [20] by $15.4$% and CLIP-S4 [23] by $13.4$% mIoU. Extensive ablation studies verify the effectiveness of each design in our framework.
@@
This paper is an extension of our conference paper [24]. To differentiate, we denote the method proposed in our conference version as ReCLIP and that in this paper as ReCLIP++. Compared to our conference version, we make further contributions, which are summarized as follows: 1) We optimize the design of the Bias Extraction Module. Specifically, we independently encode class-preference bias and space-preference bias into class-level Reference feature and patch-level positional feature respectively. Then the bias logit map, which explicitly represents both biases, is obtained by combining two features with matrix multiplication. 2) We additionally introduce a mask decoder, which takes the rectified logit map and the visual feature of CLIP as input and outputs smoother and more contextual rectified predictions. 3) We design a new strategy to generate a more accurate multi-label hypothesis for each image, which provides better supervision for the bias rectification process. 4) Benefiting from our new design, the distillation stage in ReCLIP is no longer necessary. We remove the distillation stage to simplify training and achieve even better segmentation performance. 5) We evaluate our methods on two more datasets including Cityscapes [28] and COCO Stuff [29]. On all the datasets, the segmentation performance of ReCLIP++ outperforms ReCLIP remarkably.
@@
  • We observe that when applying CLIP to pixel-level understanding tasks, unexpected biases, including space-preference bias and class-preference bias, occur. These biases may largely constrain the segmentation performance of CLIP-based segmentation models.
@@
  • We propose to explicitly encode the class-preference and space-preference bias of CLIP via learnable Reference text inputs and projection of positional embedding respectively, and model both biases into one bias logit map via matrix multiplication. Through a simple logit-subtraction mechanism and the contrastive loss built on masked features of different classes, we effectively rectify the bias of CLIP.
@@
  • We conduct extensive experiments on segmentation benchmarks under the USS setting. Experiment results show superior performance of our method to previous state-of-the-arts.
2 Related Work
@@
Pre-trained vision-language models. Pre-trained vision-language models (VLMs) [30, 31, 32, 33, 34] have developed rapidly with the help of large-scale image-text pairs available on the Internet. Recently, CLIP [19], ALIGN [35] and Slip [36] have made great progress on learning visual and textual representations jointly by using contrastive learning. With the image-level alignment with text, pre-trained VLMs have a strong ability for zero-shot classification task and can be transferred to various downstream tasks, such as object detection [37, 38] and semantic segmentation [20, 39, 40].
@@
Unsupervised semantic segmentation. While conventional approaches of semantic segmentation [7, 6] rely on pixel-level annotations and weakly-supervised methods [41, 42, 43, 44] still ask for image-level labels, unsupervised semantic segmentation (USS) methods [11, 45, 14, 16, 46] explore to train a segmentation model without any annotations. Models like [47, 48] adopt generative model [49] to separate foreground with background or generate corresponding masks. SegSort [18], HSG[12] and ACSeg [15] use clustering strategy, while IIC [10] uses mutual information maximization to perform unsupervised learning. MaskContrast [16] introduces contrastive learning into USS. Others like DSM [13] and LNE [50] exploit features extracted from self-supervised models and spectral graph theory to facilitate segmentation. However, the methods mentioned above either fail to segment images with multi-category objects or show a large performance gap with the supervised methods. Besides, they can only obtain class-agnostic masks and have to depend on additional strategies, such as Hungarian-matching algorithm [51], to match the corresponding category with the segment. Recently, pre-trained vision-language models are adopted in USS. MaskCLIP [20] modifies the image encoder of CLIP to generate patch-level features and directly performs segmentation with text features as classifiers. CLIP-py [22] performs contrastive learning between visual features from self-supervised ViT [52] and text features from CLIP. ReCo [21] performs image retrieval with CLIP and extracts class-wise embedding as classifier with co-segmentation. CLIP-S4 [23] learns pixel embeddings with pixel-segment contrastive learning and aligns such embeddings with CLIP in terms of embedding and semantic consistency. These methods can directly assign a label to each pixel, which falls into the category of language-guided unsupervised semantic segmentation. However, directly applying CLIP in USS may result in unexpected bias. No previous method considers such bias. In this paper, we propose to explicitly model and rectify the bias of CLIP for unsupervised semantic segmentation.
@@
Language-guided semantic segmentation. Recently, many works explore semantic segmentation guided by language under different settings. Zero-shot works [53, 54, 55, 56] split classes into seen and unseen sets. During the training period, only masks of seen classes are provided. For inference, models are tested on both seen and unseen classes, but the test data is still in the same distribution as the training data. Trainable open-vocabulary works [57, 58, 59, 40, 60, 61, 39, 62] are trained in one scenario with extra annotations including class-agnostic masks or image captions but are used for predicting segmentation masks of novel classes in other scenarios. Recent training-free open-vocabulary works [63, 64, 20] adapt CLIP for semantic segmentation by modifying its architecture. These methods can directly perform inference on downstream datasets without any training. From the technical view, our method also falls into the category of language-guided semantic segmentation. However, we consider the unsupervised setting. In this setting, we have access to images without any annotations during training. The training and inference images are sampled from the same distributions and the same set of categories. Such a setting is different from the typical zero-shot or trainable open-vocabulary setting.
3 Method
@@
Overview We aim to rectify the bias of CLIP including the class-preference bias and the space-preference bias, to facilitate unsupervised semantic segmentation. From a high level, class-preference bias reflects the shift of CLIP predictions towards specific classes, while space-preference bias reflects the shift of CLIP predictions towards specific spatial locations. Both biases will be finally reflected in the bias logit map. A reasonable way to rectify the bias is to subtract the bias logit map from the query logit map predicted by original CLIP.
@@
The general framework of our method is illustrated in Fig. 2. We first forward the image $I\in\mathbb{R}^{3\times H\times W}$ through the image encoder of CLIP to obtain the patch-level image feature $Z$. For each class, we manually design the text input which is named Query $Q$, and fix $Q$ throughout the training. We pass $Q$ through the text encoder of CLIP to obtain query text feature $W_{q}$ for each class, which serves as the weight of the query segmentation head. With $W_{q}$, we can obtain a query logit map $M_{q}$, which represents the segmentation ability of the original CLIP (Sec. 3.1).
@@
We extract the bias existing in CLIP in a learnable way. Specifically, we design a kind of learnable text input named Reference $R$ for each class. Passing $R$ through the text encoder of CLIP, we obtain Reference text feature $W_{r}$, which is expected to encode the class-preference bias. Meanwhile, the positional embedding is projected into the positional feature $W_{p}$ to encode the space-preference bias. Then the bias logit map $M_{b}$ is extracted via a matrix multiplication between $W_{r}$ and $W_{p}$ (Sec. 3.2).
@@
The Query logit map and the bias logit map are then passed through a Rectified Mask Generation module to generate rectified semantic masks. Firstly, a rectified logit map $M$ is generated via a simple subtraction operation between $M_{q}$ and $M_{b}$. Then, a mask decoder takes $M$ and visual feature $Z$ as input and outputs smoother and more contextual semantic masks with the help of Gumbel-Softmax operation (Sec. 3.3).
@@
To enable a more meaningful and effective bias rectification, we generate a multi-label hypothesis for each image and impose contrastive loss based on masked visual features and Query text features of different classes (Sec. 3.4). CLIP is kept frozen during the training process.
@@
Figure 2: Method overview of ReCLIP++. We propose a new framework for language-guided unsupervised semantic segmentation. We aim to rectify the class-preference and space-preference bias of CLIP via specifically designed Bias Extraction module, Rectified Mask Generation module, and Contrastive Loss module.
3.1 Baseline: Directly Segment with CLIP